Circular No. 8275

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

RADIO TRANSIENT IN SCORPIUS M. P. Rupen, A. J. Mioduszewski, and V. Dhawan, National Radio

Astronomy Observatory, report strong, on-going radio flares from the microquasar candidate associated with GSC 7861.1088 = 1RXS J162848.1– 415241 (cf. IAUC 7968). Observations with the Very Large Array (VLA) on Jan. 6, 11, and 22 UT gave flux densities at 8.46 GHz of 0.35 ± 0.08, 13.8 ± 0.8, and 6.6 ± 0.4 mJy, respectively. Quasi-simultaneous observations at 4.86 GHz on the last day gave a flux density of 4.9 ± 0.3 mJy, implying that the flux density goes as frequency to the power 0.54 ± 0.16. Such an inverted radio spectrum is almost invariably associated with the early, optically thick stage of a strong radio outburst, and this — together with the changing radio flux density — suggests that the source is currently undergoing a series of radio flares. The most reliable radio position is that from the Jan. 22 observations at 4.86 GHz: α = 16^{h}28^{m}47^{s}.283 ± 0^{s}.003, δ = −41^{o}52′39′′.05 ± 0′′.05 (equinox 2000.0), where the error bars are statistical, based on a fit using a point source plus a planar background. This is in excellent agreement with the UCAC2 optical position (Zacharias et al. 2003, from the Vizier On-line Data Catalog), which gives position end figures 47^{s}.286, 39′′.01 (quoted accuracy 0′′.01). The VLA data show no obvious signs of extension, with the best limits coming from the Jan. 11 observations, which had a beam size of 3′′.5 × 0′′.6, with the long axis oriented toward p.a. 14^{o}. A Gaussian fit to those data gave a nominal upper limit to the extent of 0′′.3 (full-width at half-maximum). Further VLA observations are planned, and observations at other wavelengths (particularly radio, x-ray, and optical) are urgently requested.

SUPERNOVA 2004I IN NGC 1072 C. L. Gerardy, and G. H. Marion, University of Texas at Austin, re-

port that a low-resolution optical spectrogram (resolution 600; range 430– 700 nm) of SN 2004I (cf. IAUC 8272), obtained on Jan. 22.13 UT with the 2.7-m Harlan J. Smith Telescope (+ LCS spectrograph) at McDonald Observatory, shows it to be a type-II supernova. The spectrum resembles that of SN 1992H, 20 days after explosion (Filippenko 1997, ARAA 35, 309), with weak Hβ and He I 587.6-nm absorption features superimposed on a blue continuum, and a strong, broad Hα emission feature with little or no blue-shifted absorption component. Adopting the 8018 km/s redshift of NGC 1072 (from the NED), the expansion velocity of the blue edge of the Hα feature is ∼ 8600 km/s.

2004 January 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green